OpenStudy (anonymous):

Heeeeelp! D: If 63.8 grams of aluminum metal (Al) react with 72.3 grams of sulfur (S) in a synthesis reaction, how many grams of the excess reactant will be left over when the reaction is complete? Be sure to write out the balanced equation for this reaction and to show all of your work.

OpenStudy (anonymous):

@dan815 @texaschic101

OpenStudy (anonymous):

I can get all the way to the part where I have to multiply Ay by the ratio, which should be 2:3, right? But when I looked this up, someone was doing it by 3/2.. I got lost there.

OpenStudy (anonymous):

@uri @Compassionate @skullpatrol

OpenStudy (anonymous):

@robtobey @johnweldon1993 Pleeeeeease.

OpenStudy (somy):

Step 1 Find moles of both

OpenStudy (somy):

you are given masses so mole= mass/ molecular mass and get mole of each

OpenStudy (somy):

then choose which ever mole you want (i'd suggest you to choose the one that has smaller value ~_~ well its usually faster for me) now look at the ratio between Al and S in your balanced reaction and now using mole of one of them and ratio of both of them -find a new mole

OpenStudy (somy):

if the new mole is LESS than mole that was originally given for that element in the question = that means this element is in Excess

OpenStudy (somy):

which means you found the 'required mole' so subtract mole that you found using ration from mole that was given in the question

OpenStudy (somy):

resulting mole will be that Excess not used mole so now u can find its mass mole= mass/ molecular mass mass= mole* molecular mass = mole that you got by subtraction * molecular mass of that element and get your final answer :)

OpenStudy (anonymous):

So.. Okay. I get the process, but where you multiply one of the moles by the ration.. You can choose either one? And do you multiply it by 2/3, or 3/2?

OpenStudy (somy):

yeah you can choose whichever you want because when you take one of them and if its in excess- you would get MORE mole of the other reactant meaning more than its given in the question (i hope u r not confused, when i say you take mole of something - i mean for the ratio to find mole of the other thing- so if mole of other thing you will get LESS than its given for it in the question= that means its in Excess= which is what we want, and so this Less mole is the mole required for the reaction and the rest is Excess that is not used)

OpenStudy (somy):

write a balanced reaction so that i can tell what are you going to do

OpenStudy (anonymous):

2Al+3S-->Al2S3 Right @Somy ?

OpenStudy (jfraser):

your balanced reaction is right, now find the available moles of each to find out which one will run out first

OpenStudy (somy):

yeah good so now take whichever mole you want if you take Al mole then it'll be like this ratio is 2:3 so Al mole - 2 X mole - 3 X mole= Al mole * 3 / 2 = X mole of S REQUIRED if you take S mole ratio 2:3 so S mole - 3 Y mole - 2 Y mole= S mole * 2 /3 = Y mole of Al REQUIRED so now when you get REQUIRED for Al and S, compare them with the amount given in the question if one of these values of Al and S is LESS then amount given in the question that means that one is in Excess

OpenStudy (anonymous):

2 Al + 3 S → Al2S3 (63.8 g Al) / (26.98154 g Al/mol) = 2.3646 mol Al (72.3 g S) / (32.0655 g S/mol) = 2.2548 mol S 2.2548 moles of S would react completely with 2.2548 x (2/3) = 1.5032 moles of Al, but there is more Al present than that, so Al is in excess. ((2.3646 mol Al initially) - (1.5032 mol Al reacted)) x (26.98154 g Al/mol) = 23.2 g Al left over

OpenStudy (somy):

@Tech1 good job you did there :) but from now on please leave solving part for the asker :)